# Cross-Subject DD: A Cross-Subject Brain-Computer Interface Algorithm

### v1

@@ 4.3.1 Decoding performance across different models
Table LABEL:table_1 comprehensively presents the performance of the proposed cross-subject decoding strategy compared with other baseline models on the BCIC IV 2a dataset. It includes the classification accuracy for each subject and the average accuracy comparison. The proposed CSDD model demonstrates excellent performance in cross-subject decoding tasks, achieving the highest average classification accuracy and significantly surpassing other models.
Firstly, in terms of average classification accuracy, the CSDD model outperforms EEGNet by 13.53%, with a statistically significant difference ($\mathit{p}$ < 0.01). This indicates that the CSDD model exhibits stronger robustness and generalization ability when extracting common features and decoding across different subjects. Additionally, compared to CTNet, the CSDD model achieved a 6.21% higher average classification accuracy and a higher kappa value, while also demonstrating a lower standard deviation ($\mathit{p}$ < 0.05). When compared with other baseline models, such as the Multi-branch 3D CNN, CCNN, and CRAM, the CSDD model achieved classification accuracy improvements of 12.68%, 9.23%, and 5.63%, respectively, with these differences also being statistically significant ($\mathit{p}$ < 0.01). These results further validate the superiority of the CSDD model in cross-subject tasks, particularly in its effective utilization of common features during feature extraction and model training to reduce the impact of individual differences on decoding performance. Although the CSDD model showed a 3.28% improvement in accuracy compared with ShallowConvNet, this difference was not statistically significant ($\mathit{p}$ > 0.05). This suggests that these models may perform similarly in some cases, which may be attributed to the limitations of these baseline models in handling inter-subject differences. This further highlights the advantages of the CSDD model in terms of adaptability and generalization.
@@
4.3.2 Effect of common feature extraction on cross-subject decoding performance
This section investigates the impact of varying predefined significance thresholds for common features extraction on the cross-subject decoding performance of the CSDD model, using the BCIC IV 2a dataset. The common features were extracted at different predefined thresholds $\alpha_{\text{threshold}}$ ( 0.1, 0.05, 0.01, and 0.001 ), and the performance of the model was evaluated.
The classification accuracy results based on these thresholds are summarized in Fig. 8. When $\mathit{p}$ < 0.1, the average accuracy was 62.96%, with subject-specific accuracies ranging from 49.31% (Subject-2) to 65.62% (Subject-1). Performance improved as the $\mathit{p}$-value threshold was reduced: when $\mathit{p}$ < 0.05, accuracy increased to 63.58%, and when $\mathit{p}$ < 0.01, it rose to 64.93%, with Subject-4, Subject-6, and Subject-8 showing substantial improvement. However, when $\mathit{p}$ < 0.001, accuracy reached 63.79%, showing a marginal improvement over when $\mathit{p}$ < 0.01, indicating that further reduction in features may not provide significant performance gains.
@@ 4.3.2 Effect of common feature extraction on cross-subject decoding performance
Figure 9: The visualization for Subject-3. (a) T-SNE visualization of raw data. (b) T-SNE visualization of the feature-extracted data. Different colors represent different categories.
@@ 5 Discussion
In this study, we proposed the CSDD model, which leverages common feature extraction to improve cross-subject decoding performance in BCI tasks. The CSDD model demonstrated superior decoding results across subjects, significantly outperforming baseline models, particularly in terms of classification accuracy and generalization ability. The results show that the CSDD model effectively handles inter-subject variability by extracting shared features, leading to improved decoding performance across subjects.
This paper involves three research directions. The first is the study of cross-subject universal BCI. The second is the method for distinguishing between common and subject-specific features. The third is the system filter.
5.1 Cross-subject universal BCI decoding model
In cross-subject BCI research, traditional BCI models are mainly based on personalized training strategies, where models are individually trained for specific brain activity patterns of each subject[36][37]. Although this approach performs well in individual tasks, it has significant limitations. Personalized training typically requires large amounts of labeled data and extensive training time, which becomes computationally expensive and difficult to scale, especially when dealing with a large number of subjects. Additionally, individual differences introduce biases in each subject’s model training, resulting in poor generalization across subjects. In recent years, with advancements in technology, methods such as transfer learning and domain adaptation have been proposed to reduce the need for personalized training and improve cross-subject adaptability[38][39][40]. However, these methods do not yet fully address the problem of poor adaptability, primarily because they rely more on calibrating individual differences and lack an effective mechanism to extract common features across subjects.
Unlike traditional personalized BCI models, the cross-subject universal BCI model (CSDD) proposed in this paper introduces the concept of extracting common features across subjects. The innovation of this model lies in its reliance on common features shared across subjects rather than subject-specific features, constructing a universal model. This approach not only reduces the dependence on personalized data but also significantly improves the adaptability and decoding accuracy for new subjects, thus achieving better generalization across subjects.
Fig. 9 uses t-SNE (t-distributed Stochastic Neighbor Embedding) [41] to reduce the dimensionality of the data from subject 3 and displays the distribution of extracted common features across different subjects. The experimental results show that the t-SNE method successfully separates the four motor intentions (left hand, right hand, foot, and tongue), with clear separation between feature distributions of different classes. Notably, the cross-subject accuracy for Subject-3 was 80.38%, emphasizing the effective feature clustering for different motor intentions in this subject. The features extracted for subject 3 are well-clustered in the feature space, indicating that the proposed method can effectively extract common features across subjects, thereby enhancing the generalization ability of the model and achieving accurate decoding. In contrast, raw EEG signals without processing show a significant overlap in the feature space, making classification more difficult. This phenomenon further validates the robustness of the proposed method, demonstrating its ability to extract easily classifiable features from complex and noisy EEG signals, significantly improving cross-subject decoding accuracy.
5.2 Distinguish between common and subject-specific features
Traditional BCI decoding methods typically rely on optimizing the temporal and spatial dimensions or adjusting the decoding functions to enhance the learning of common features across subjects[42][43][44]. Deep learning methods, on the other hand, approximate these common features through large-scale data training, thereby improving decoding accuracy[45][46][47]. Although these methods improve model generalization by increasing common features across individuals and reducing subject-specific features, they typically rely on a single strategy and lack clear methods for distinguishing and optimizing common and subject-specific features. In contrast, the DD algorithm, as a white-box model, offers lower computational complexity and strong generalization capability, capable of expressing both the independent and interactive effects of input variables on output through the relation spectrum.
Therefore, this paper proposes DD Filter, which can simplify the formula of the trained personalized model into a polynomial form, using random probability statistical tests to select statistically significant common relation items and filter out subject-specific relation items.
The significance level is controlled through one sample $t$-test, which regulates the number of common relation items. As shown in Fig. 7, as the predefined threshold $\alpha_{\text{threshold}}$ decreases from 0.1 to 0.001, the number of extracted common relation items progressively decreases. This indicates that more stringent significance levels select fewer features. In Fig. 8, decoding accuracy fluctuates with variations in the $\mathit{p}$-value. Although a slight improvement is observed when $\mathit{p}$ < 0.1, the accuracy change is minimal and does not lead to a significant performance enhancement under stricter thresholds. This suggests that under excessively strict significance levels (e.g., $\mathit{p}$ < 0.001), although the number of features is reduced, their contribution to cross-subject decoding is limited and may even detract from the performance of the model. This phenomenon further underscores the necessity of finding an appropriate balance between feature quantity and decoding performance during model optimization. Proper feature selection and avoidance of overfitting are essential for effectively enhancing cross-subject decoding performance.
5.3 System component filtering method with a filter-like structure
Traditional filtering methods, especially in signal processing, typically rely on Fourier transforms to convert time-domain signals into the frequency domain. This allows for the removal of noise while retaining useful frequency components. However, despite its success in signal level applications, this concept has not yet been effectively applied to filtering out irrelevant components at the system level.
The main contribution of this paper is the extension of the filtering concept to relation system processing. Specifically, a relation filtering mechanism is introduced to identify and remove irrelevant or redundant relation components within a system.
In the proposed framework, the personalized model for each subject is initially transformed into a relation spectrum, similar to the transformation of time-domain signals into frequency spectra in traditional signal processing. This transformation allows subject-specific EEG decoding features to be represented within a unified relation spectrum, making it easier to compare and analyze patterns across subjects. By applying statistical analysis, shared components across individuals are identified, while individual differences are treated as “noise” and filtered out.
Unlike conventional time-frequency feature filtering methods, the proposed approach focuses on filtering at the relational level of the system rather than merely extracting time frequency features. By capturing shared relational structures across subjects and eliminating redundant information caused by individual variability, the proposed filtering method aims to optimize the cross-subject representation of EEG data. The goal is to remove unnecessary subject-specific information and retain only those common features that generalize well across subjects, thus enabling the construction of a more universal and adaptive BCI model.
@@ 6 Conclusion
In summary, this study proposes a novel cross-subject BCI decoding model based on the extraction of common features. By training personalized models for each subject and subsequently extracting common relation items across subjects through the system filter proposed in this study, a generalized cross-subject BCI model has been successfully constructed, achieving complete cross-subject decoding. Experimental results demonstrate that the proposed method effectively improves decoding accuracy for new subjects, offering better generalization compared to traditional methods. Under cross-subject conditions, the method achieved an average accuracy of 64.85% on the BCIC IV 2a dataset.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (62303423), the STL 2030-Major Project (2022ZD0208500), Postdoctoral Science Foundation of China (2024T170844, 2023M733245), the Henan Province key resaerch and development and promotion of special projects (242102311239), Shaanxi Provincial Key Research and Development Program (2023GXLH-012).
Data availability
This study utilizes publicly available datasets, with the corresponding data links provided in the manuscript.

### v2

@@ 4.3.1 Decoding Performance Across Different Models
Table 3 comprehensively presents the performance of the proposed CSD-SF compared with other baseline models on the BCIC IV 2a dataset.
Firstly, in terms of average classification accuracy, the proposed CSD-SF outperformed EEGNet by 13.53%, with a statistically significant difference ($\mathit{p}<0.01$). This demonstrates the superior robustness and generalization ability of CSD-SF in extracting common components and enabling cross-subject decoding. Additionally, compared with CTNet, CSD-SF achieved a 6.21% higher average accuracy and a higher kappa value, while also demonstrating a lower standard deviation ($\mathit{p}<0.05$), indicating more stable performance across subjects. When compared with other baseline models, such as the Multi-branch 3D CNN, CCNN, and CRAM, CSD-SF achieved classification accuracy improvements of 12.68%, 9.23%, and 5.63%, respectively, with these differences also being statistically significant ($\mathit{p}<0.01$). These results further validate the superiority of CSD-SF in cross-subject tasks, particularly in its effective utilization of common components during feature extraction and model training to reduce the impact of individual differences on decoding performance. Although CSD-SF achieved a 3.28% improvement over ShallowConvNet, the difference was not statistically significant ($\mathit{p}>0.05$). This suggests that these models may perform similarly in some cases, which may be attributed to the limitations of these baseline models in handling inter-subject differences. This further highlights the advantages of CSD-SF in terms of adaptability and generalization.
@@
4.3.2 Effect of Common Components Extraction on Cross-Subject Decoding Performance
The classification accuracy results based on these thresholds are summarized in Fig. 9. When $\mathit{p}<0.1$, the average accuracy was 62.96%, with subject-specific accuracies ranging from 49.31% (Subject-2) to 65.62% (Subject-1). Performance improved as the $\mathit{p}$-value threshold was reduced. When $\mathit{p}<0.05$, accuracy increased to 63.58%, and when $\mathit{p}<0.01$, it rose to 64.93%, with Subject-4, Subject-6, and Subject-8 showing substantial improvement. However, when $\mathit{p}<0.001$, accuracy reached 63.79%, showing a marginal improvement over when $\mathit{p}<0.01$, indicating that further reduction in features may not provide significant performance gains.
@@ 4.3.2 Effect of Common Components Extraction on Cross-Subject Decoding Performance
Figure 10: Cross-subject decoding accuracy with and without the system filter.
Figure 11: The visualization for Subject-3. (a) T-SNE visualization of raw data. (b) T-SNE visualization of data without using the system filter. (c) T-SNE visualization of data using the system filter (CSD-SF). Different colors represent different categories.
4.4 Ablation Study
To evaluate the effectiveness of the proposed system filter in the CSD-SF model for improving cross-subject EEG decoding, an ablation study was performed. Specifically, we compared two conditions for cross-subject decoding using the LOSO cross-validation: with system filter (CSD-SF) and without system filter. In the “without system filter” condition, no relation spectrum transformation or filtering was applied. Instead, cross-subject decoding was performed directly using deep feature extraction followed by a fully connected layer. The experimental results are illustrated in Fig. 10. Our proposed method achieved high classification accuracy across all subject, with significant improvements observed for Subject 3 and Subject 5. On average, the classification accuracy increased by 8.62% compared to the “without system filter” condition ($\mathit{p}<0.05$). These results demonstrate that system filter achieved superior performance in the four-classification MI task, underscoring its effectiveness in enhancing cross-subject generalization.
The t-SNE(t-distributed Stochastic Neighbor Embedding)[38] visualizations for Subject 3 are shown in Fig. 11, highlighting the impact of the system filter MI task separation. The raw EEG data exhibits significant overlap between the different motor imagery tasks (Left Hand, Right Hand, Foot, and Tongue), suggesting that individual-specific noise impedes the separation of tasks and complicates classification. In the condition without the system filter, the leave-one-subject-out cross-validation was performed on the extracted features without applying any transformation or component filtering. While some task separation is observed, the clusters still overlap significantly, indicating that subject-specific noise continues to affect model performance. In contrast, after applying the system filter (CSD-SF), the separation between the MI tasks becomes much clearer. By focusing on common components shared across subjects and filtering out personalized components, the system filter enhances the clarity of the clusters, improving task separation and decoding accuracy. These results demonstrate that the system filter significantly improves the cross-subject decoding performance by extracting stable, generalized features and reducing the impact of individual differences.
@@ 5 Discussion
In this paper, we propose a system filter that extends the traditional concept of signal filtering to the system modeling level and apply it to cross-subject MI-EEG decoding, termed CSD-SF. The system filter enables explicit decomposition of system components, filtering out personalized components and retaining common components across subjects to construct a universal decoding model. Experimental results demonstrate that CSD-SF achieves superior performance in the four-class MI task, confirming its effectiveness in improving model generalization.
This discussion highlights two key aspects: the methodological value of the system filter as a general modeling framework, and the role of the CSD-SF in developing universal cross-subject decoding models based on common components.
5.1 System Filter as a General Modeling Framework
The proposed system filter extends the notion of filtering from the signal domain to the system domain. Traditional signal filtering, such as Fourier or wavelet filtering, aims to remove noise while preserving the essential components of the signal. The system filter extends this concept by operating directly on the system representation, focusing on the intrinsic relationships among variables rather than the raw signal itself. By expanding the system into a polynomial representation and mapping it into the spectral domain, it enables explicit decomposition of relational components, allowing unnecessary items to be selectively filtered out.
This structural filtering provides a controllable and interpretable approach to system modeling. Rather than relying on implicit parameter adaptation or end-to-end optimization, it imposes a structural stability constraint to preserve essential mappings while filtering out unstable relations[39, 40, 41, 42, 43]. By explicitly constraining system structure and selectively removing unwanted components, the system filter achieves interpretable and generalizable modeling. Although this paper focuses on EEG decoding, the same principle can be extended to eliminate unwanted components in other complex systems, offering a general paradigm for stable and interpretable modeling.
5.2 Toward Universal Cross-Subject BCI Decoding Models
Traditional approaches, including temporal–spatial filter optimization [44, 45, 46] and deep learning models trained on large datasets [47, 48, 49], achieve adaptability by reinforcing cross-subject common features or accommodating subject-specific variations. However, they do not explicitly distinguish common components from personalized ones, and residual individual interference often limits their generalization ability.
In contrast, the proposed CSD-SF framework, built upon the system filter, provides an explicit and interpretable mechanism for isolating common and personalized components in cross-subject EEG decoding. The method expands individual system models into relation spectrums and applies one-sample $t$-test to evaluate the stability of each relation item. Unstable items are regarded as personalized components reflecting individual variability, whereas stable relation items consistent across subjects are retained to construct the decoding model. These stable relation items are assumed to represent shared cortical activation patterns underlying motor imagery.
By explicitly distinguishing common from personalized relation items at the system level, CSD-SF suppresses individual noise while preserving common neural dynamics. As shown in Fig. 11(c), the common model derived from stable relation items achieves clear task-specific separation when decoding new subjects, confirming the effectiveness of the proposed method.
Beyond the methodological contributions, the CSD-SF framework enhances the practical usability of EEG-based BCI systems. By identifying neural patterns that remain stable across individuals, it reduces the need for extensive subject-specific calibration and enables faster adaptation in clinical and rehabilitation contexts. This advantage is critical for motor-impaired users, for whom minimizing recalibration demands is essential to ensure consistent and reliable BCI performance.
@@ 6 Conclusion
This paper proposes a novel system filter, a method that transforms a complex system into a series of polynomial terms. By expanding the system into a spectral representation, the filter selectively removes unnecessary components, retaining only the target components necessary for the task.
In the context of cross-subject decoding, we applied the system filter within the CSD-SF framework, specifically developed for MI-EEG decoding. The personalized models were transformed into relation spectrums, and statistical testing across subjects enabled the separation of personalized components from common components. The retained common components were used to construct a universal model.
Experimental results on the BCIC IV 2a dataset demonstrate the effectiveness of the system filter in improving cross-subject decoding performance. The proposed method achieved an average accuracy improvement of 3.28% over baseline approaches, underscoring the significance of focusing on common components to enhance cross-subject generalization. This improvement confirms that the system filter enhances the robustness of the decoding model by isolating common components and removing personalized noise, thereby improving cross-subject generalization. This is crucial for enabling effective BCI applications across diverse subjects.